Solve (v - 8)^2 - 54 = 0, where v is a real number.
Simplify your answer as much as possible.
If there is more than one solution, separate them with commas.

Answer by stanbon(75887)

Solve (v - 8)^2 - 54 = 0
Factor:
[(v-8)-sqrt(54)][(v-8)+sqrt(54)] = 0
---
v-8 = sqrt(54) or v-8 = -sqrt(54)
--
v = 8+sqrt(54) or v = 8-sqrt(54)
==================
Cheers,
Stan H.